### v1

@@ 4.4 Temporal Transformer
The multimodal tokenizer provides RGB and pose tokens for each micro-action. Since an action segment consists of a sequence of micro-actions, these micro-action tokens are aggregated over time via a temporal transformer. A video $\mathbf{V}$, divided into $K$ micro-actions can be represented by two sets of tokens $\{\hat{f}_{k}^{\text{RGB}}\}_{k=1}^{K}$ and $\{\hat{f}_{k}^{\text{Pose}}\}_{k=1}^{K}$, respectively, produced by the multimodal tokenizer. These $2\times K$ tokens of dimension $d$ form the input sequence of the temporal transformer. Positional encoding and modality embedding are added to each input token to indicate the temporal location and source modality. We use the fixed sine/cosine positional encoding [60], where we assign the same position to two tokens coming from the same micro-action. Modality embeddings are learned and shared across the tokens from the same modality (RGB or Pose). Following standard practice [17, 18], we prepend an additional learnable class token [CLS] $\in\mathbb{R}^{d}$, the output of which is then used for classifying actions.
@@ 4.5 Learning Objectives
HandFormer is trained end-to-end for action recognition, supervised via a cross-entropy loss: $\mathcal{L}_{cls}=-a_{i}\log\hat{a}_{i}$, where $a_{i}$ represents the ground truth action label for the $i^{th}$ sample, and $\hat{a}_{i}$ denotes the predicted action category by HandFormer. The pose and RGB modalities, serving as inputs, provide complementary information regarding a scene, capturing motion and interacting objects, respectively. To effectively utilize this information, we employ explicit verb and object supervision ($\mathcal{L}_{verb}$, $\mathcal{L}_{obj}$) through two additional learnable class tokens. Given that pose strongly correlates with verb class, the verb class token selectively attends to pose encodings. Similarly, as objects are reliably identifiable from RGB frames alone, the object class token exclusively attends to frame features.
@@
[…]
where $\Phi_{\text{ant}}$ denotes a linear projection layer.
Hence, the total loss is:
@@ 5 Experiments
Datasets. We conduct experiments on two publicly available hand-object interaction datasets with multiple views, Assembly101 [51] and H2O [33]. Assembly101 [51] is a large-scale multiview dataset that features videos of procedural activities for assembling and disassembling 101 take-apart toy vehicles. This dataset has 1380 fine-grained actions resulting from a combination of 24 verbs and 90 objects. Additionally, 3D hand pose estimations are provided. The dataset consists of 12 temporally synchronized views — 8 static and 4 egocentric. The egocentric views exhibit low-resolution monochrome images, making it challenging even for human eyes to discern objects in view. As a result, we opt for fixed views in our RGB modality. H2O [33] features four participants performing 36 actions that involve 8 different objects and 11 verbs. 3D hand poses with 21 key points per hand are provided, along with 6D object poses, camera poses, object meshes, and scene point clouds. It consists of 5 temporally synchronized RGB views — 4 fixed and 1 egocentric. In our work, we use hand poses and RGB frames from the egocentric view only.
Implementation Details. Similar to [66], our pose input consists of 120 frames by setting $T^{\prime}=120$. The number of frames per micro-action ($N$) is 15. For multimodal experiments, we set the window stride $R=N$ and take $K=8$ non-overlapping micro-actions, which allows us to sample 8 RGB frames for each video following [51]. However, we allow a 50% overlap between consecutive micro-actions for pose-only versions. Unless otherwise specified, we utilize the pre-trained ViT-g/14 and ViT-L/14 models from DINOv2 [46] followed by a linear layer without any fine-tuning as the frame encoder $F$ for Assembly101 [51] and H2O [33], respectively. We also evaluate a lightweight alternative by fine-tuning ResNet-50 [28] in the Supplementary. Each model is trained for 50 epochs using SGD with momentum 0.9, a batch size of 32, a learning rate of 0.025, and step LR decay with a factor of 0.1 at epochs {25, 40}. During training, the frame encoder $F$ is kept frozen except for the final linear layer.
Model Variants. We propose several variants of our model by adjusting the width $d$ and the number of layers $T_{n}$ of the transformer to strike a balance between efficiency and accuracy. Our default HandFormer, denoted HandFormer-B, has parameters $(d,T_{n})=(256,2)$. We introduce a larger variant, HandFormer-L, with $(d,T_{n})=(512,4)$. We also explore different configurations for the number of input joints $J$ per hand. Unless otherwise mentioned, we utilize all 21 joints per hand along with the base model denoted as HandFormer-B/21 while offering a highly efficient option utilizing only 6 joints per hand (5 fingertips and the wrist) termed HandFormer-B/6.
@@ 5.1 Comparison with State-of-the-Art
To evaluate the effectiveness of our proposed architecture, we compare it against several baselines. For pose-only comparisons, we choose a graph-based network MS-G3D [43] and an attention-based method ISTA-Net [66] — the two best performing skeleton-based methods on Assembly101 [51], as reported by [66]. We also employ video baselines TSM [39] and SlowFast [23], which emphasize efficiency and high temporal resolution, respectively. Additionally, we include Cho et al. [13], the state-of-the-art for the H2O dataset. For Assembly101, we replicate the results of the aforementioned methods using RGB frames from view 4 of the dataset, while results on H2O are acquired from the respective papers. Furthermore, on both datasets, we train and test RGBPoseConv3D [20], state-of-the-art in multi-modal action recognition with skeleton and RGB data.
We evaluate three variants of our method by controlling the input modalities. As shown in Tab. 1, our unimodal pose-only model excels in verb recognition on Assembly101. In contrast, RGB-based methods benefit from object appearance and usually perform well in terms of action accuracy, which can be attributed to their high object recognition performance. In this context, the accuracy of ISTA-Net on H2O is not directly comparable to our method, as they also use the 6D object poses, while we only use hand poses as input. RGBPoseConv3D struggles to achieve satisfactory performance, particularly on Assembly101, suggesting that generalizing to hand poses is non-trivial for skeleton-based methods. Therefore, we combine two best-performing unimodal methods with late fusion (MS-G3D + TSM) to create a stronger baseline. Our model even outperforms this, indicating the effectiveness of our proposed multimodal fusion.
Method Pose RGB Assembly101 H2O Action Verb Object Action MS-G3D [43] 28.78 63.46 37.26  50.83 ISTA-Net [66] 28.14 62.70 36.77  89.09 SlowFast [23] - - -  77.69 TSM [39] 35.27 58.27 47.45  - Cho et al. [13] - - -  90.90 RGBPoseConv3D [20] 33.61 61.99 42.90  83.47 MS-G3D + TSM 39.74 65.12 51.12  - HandFormer-B/21 28.80 65.33 36.28  57.44 32.07 55.61 44.89  84.71 41.06 69.23 51.17 93.39
Table 1: Quantitative comparison with state-of-the-art methods on Assembly101 [51] and H2O [33]. For H2O, 6D object pose is used by ISTA-Net during training and inference and by Cho et al. during training. ‘MS-G3D + TSM’ denotes a late fusion of the corresponding unimodal architectures.
@@ 5.2 Skeleton-based Action Recognition for Hands
The compositional nature of action classes in hand-object interaction videos allows us to break down the action into a verb and an object. Recognizing such actions from 3D hand poses is an ill-posed problem, as the hand skeletons lack explicit information about the interacting objects, which are also a part of the action semantics. However, as the pose data completely captures the hand motion information, it can be reliably used for verb recognition. Therefore, we evaluate a pose-only version of our method on the verb recognition task and compare the performance and efficiency metrics with other state-of-the-art skeleton-based methods in Fig. 5. Our method uses significantly fewer GFLOPs due to our spatiotemporal factorization using micro-actions. With $J=21$, all our HandFormer variants outperform existing methods, while $J=6$ variant is extremely efficient with comparable accuracy. HandFormer-H/21 is a combined variant of HandFormer-B and HandFormer-L with $J=21$, which is our best-performing model, improving over MS-G3D by 2.04% while maintaining comparable FLOPs.
@@ 5.3 How many RGB frames are required?
In our model, the pose modality maintains a high temporal resolution to capture fine-grained hand movements, resulting in good verb recognition performance. On the contrary, RGB frames are primarily required to introduce semantic context beneficial for object recognition and do not necessitate a high temporal resolution like hand movements. The design of our model allows us to sample only a few RGB frames (as low as one) but still perform competitively at a reduced computational cost.  Fig. 6 shows the impact of using more RGB frames for Assembly101 [51]. Using only one RGB frame in HandFormer (35.46) outperforms the video model TSM (35.27), as shown in Tab. 1. This performance gain stems primarily from a significant improvement in object accuracy, with only a slight enhancement in verb accuracy. However, including more RGB frames shows a diminishing return, which is unsurprising as additional frames are expected to provide redundant information. These results are obtained with a simplified version of our model by setting $\mathcal{L}=\mathcal{L}_{cls}$ and bypassing the multimodal tokenizer.
@@
[Uncaptioned image] Figure 5: Comparison of skeleton-based methods for verb recognition on Assembly101 [51]. Our method achieves state-of-the-art performance while utilizing significantly fewer FLOPs.
@@
[Uncaptioned image] Figure 6: Ablating number of RGB frames in HandFormer on Assembly101 [51]. With more RGB frames, verb recognition shows marginal gain, whereas object recognition shows improvement with diminishing returns.
5.4 Are 3D hand poses an efficient alternative for multi-view?
@@
Multi-view action recognition, while benefitting from precise hand-movement information in 3D space, processing all views with video models is expensive and highly redundant. In Tab. 2, we demonstrate that combining 3D hand pose with a single RGB view (view 4) achieves comparable performance to multi-view action recognition on Assembly101 [51]. Specifically, our action recognition performance (‘Single View + RGB’) matches the fusion of the two most informative views — view 1 and view 4. Notably, our verb recognition accuracy is similar to the combination of all 8 RGB views while using 3×3$\times$3 × fewer FLOPs (see Supplementary). Additionally, using hand pose in combination with an RGB view (‘Single View + RGB’) outperforms directly using the egocentric videos (‘Single View + Egocentric’) from which the hand poses are derived. While fusing multiple RGB views improves accuracy by ensembling multiple complementary predictions, the computational overhead also increases significantly. In contrast, our model processes hand pose and single-view RGB frames, enhancing efficiency by leveraging the less redundant and low-dimensional pose data.
@@
Views Action Verb Object Single View 35.27 58.27 47.45 Single View + Egocentric 37.75 61.80 49.43 Two Views 41.96 65.22 53.26 All (8) Views 47.51 70.99 57.73 Single View + Pose 41.06 69.23 51.17 Table 2: Multi-view Action Recognition on Assembly101[51]. HandFormer with ‘Single View+Pose’ has better performance than ‘Single View+Egocentric’. However, verb performance is comparable to using all (8) RGB views. Method Trained on Tested on Assembly101[51] Action Verb Object TSM [39] v4 + ego v4 + ego 37.75 61.80 49.43 v4 + ego v1 + ego 35.27 59.53 47.72 v1 + ego v1 + ego 36.21 60.78 48.52 Our Method v4 + Pose v4 + Pose 41.06 69.23 51.17 v4 + Pose v1 + Pose 38.43 67.86 48.32 Table 3: Cross-view performance of HandFormer shows its generalization capability to unseen view 1, outperforming the video baseline which is trained on view 1 directly. Egocentric views are the source of hand poses in Assembly101 and, therefore, are included in the video models.
@@
Cross-view Generalization. 3D hand pose offers a unique opportunity for cross-view generalization because of its universality across different viewpoints. To evaluate the effectiveness of our method for unseen views, we train our model with frame-wise RGB features from view 4 and test it on view 1. As a baseline video model, we train TSM [39] on both RGB views separately. As our method includes a 3D hand pose, which is obtained from egocentric views, we include the egocentric videos in the TSM baselines for a fair comparison. As seen from Tab. 3, our method, trained on view 4, generalizes well on unseen view 1, outperforming the TSM model that was directly trained on view 1.
@@ 5.5 Egocentric Action Recognition
Recognizing actions in egocentric videos is challenging due to factors such as camera motion and the occlusion of interacting objects by hand. Additionally, in the case of Assembly101 [51], the egocentric cameras are similar to Oculus Quest VR headsets, which provide monochrome low-resolution frames. As a result, action recognition performance is significantly lower compared to fixed RGB views, as demonstrated in [51]. We address this challenging scenario with our proposed multi-modal architecture and achieve state-of-the-art performance in egocentric action recognition on Assembly101. For this experiment, we used frame-wise TSM features provided by [51]. As depicted in Tab. 4, our model, using a single egocentric view (e3 or e4) outperforms the fusion of four egocentric views as reported in [51]. Moreover, fusing e3 and e4 significantly enhances our model, resulting in a 4.25% increase in action accuracy over the baseline.

### v2

@@ 4.4 Temporal Transformer
The multimodal tokenizer provides RGB and pose tokens for each micro-action. Since an action segment consists of a sequence of micro-actions, these micro-action tokens are aggregated over time via a temporal transformer. A video $\mathbf{V}$, divided into $K$ micro-actions can be represented by two sets of tokens $\{\hat{f}_{k}^{\text{RGB}}\}_{k=1}^{K}$ and $\{\hat{f}_{k}^{\text{Pose}}\}_{k=1}^{K}$, respectively, produced by the multimodal tokenizer. These $2\times K$ tokens of dimension $d$ form the input sequence of the temporal transformer. Positional encoding and modality embedding are added to each input token to indicate the temporal location and source modality. We use the fixed sine/cosine positional encoding [60], assigning the same position to two tokens from the same micro-action. Modality embeddings are learned and shared across the tokens from the same modality (RGB or Pose). Following standard practice [17, 18], we prepend an additional learnable class token [CLS] $\in\mathbb{R}^{d}$, the output of which is then used for classifying actions.
@@ 4.5 Learning Objectives
HandFormer is trained end-to-end for action recognition, supervised by a cross-entropy loss: $\mathcal{L}_{cls}=-a_{i}\log\hat{a}_{i}$, where $a_{i}$ represents the ground truth action label for the $i^{th}$ sample, and $\hat{a}_{i}$ denotes the predicted action category. The input pose and RGB modalities provide complementary information regarding a scene, capturing motion and interacting objects, respectively. To effectively utilize this, we employ explicit verb and object supervision ($\mathcal{L}_{verb}$, $\mathcal{L}_{obj}$) via two additional learnable class tokens. Given that pose strongly correlates with verbs, whereas objects are identifiable from RGB frames alone, we let the verb class token attend exclusively to pose encodings and the object class token to frame features.
Feature Anticipation Loss. Hand pose sequence captures the primary sources of state changes during hand-object interactions. We posit that the visual state from an initial RGB frame, in combination with the subsequent hand pose sequence, is indicative of the visual state that results from the completion of the sequence. Therefore, given an RGB feature and the corresponding pose features from a micro-action, we force our model to anticipate the RGB feature for the next micro-action by minimizing an $L_{1}$ feature loss. This loss, inspired by existing efforts [62, 25], quantifies the difference between the anticipated image feature and the true feature extracted from a frozen image encoder. Formally,
@@
where $\Phi_{\text{ant}}$ denotes a linear projection layer. Hence, the total loss is-
@@ 5 Experiments
Datasets. We conduct experiments on two publicly available hand-object interaction datasets. Assembly101 [52] is a large-scale multiview dataset that features videos of procedural activities for assembling and disassembling 101 take-apart toy vehicles. This dataset has 1380 fine-grained actions based on 24 verbs and 90 objects. It consists of 12 temporally synchronized views — eight static and four egocentric. The egocentric views exhibit low-resolution monochrome images, making it challenging even for human eyes to discern objects in view. Hence, we opt for fixed views as our RGB modality. 3D hand poses estimated using off-the-shelf MEgATrack [27] are provided by the dataset. H2O [34] features four participants performing 36 actions that involve 8 objects and 11 verbs. Ground truth 3D hand poses are provided. The dataset consists of 5 temporally synchronized RGB views — 4 fixed and 1 egocentric. We use RGB frames from the egocentric view only.
Implementation Details. Similar to [66], our pose input consists of 120 frames by setting $\mathcal{T}^{\prime}=120$. The number of frames per micro-action ($N$) is 15. For multimodal experiments, we set the window stride $R=N$ and take $K=8$ non-overlapping micro-actions, which allows us to sample 8 RGB frames for each video following [52]. However, we allow a 50% overlap between consecutive micro-actions for pose-only variants. A frozen ViT [18] or ResNet [29], followed by a learnable linear layer, is used as the frame encoder $F$. More details are provided in Suppl. Sec. C. Each model is trained for 50 epochs using SGD with momentum 0.9, a batch size of 32, a learning rate of 0.025, and step LR decay with a factor of 0.1 at epochs {25, 40}. Loss hyperparameters $\{\lambda_{1},\lambda_{2},\lambda_{3}\}$ are chosen to be $\{1.0,1.0,2.0\}$.
Model Variants. We propose several variants of our model by adjusting the width $d$ and the number of layers $T_{n}$ of the transformer to balance efficiency and accuracy. Our default HandFormer, denoted HandFormer-B, has parameters $(d,T_{n})=(256,2)$. We introduce a larger variant, HandFormer-L, with $(d,T_{n})=(512,4)$. We also explore different configurations for the number of input joints $J$ per hand. Unless otherwise mentioned, we utilize all 21 joints per hand along with the base model denoted as HandFormer-B/21 while offering a highly efficient option utilizing only six joints per hand (five fingertips and the wrist) termed HandFormer-B/6. Additionally, HandFormer-X/J×$\times\mathcal{N}$ denotes a multimodal variant that utilizes $\mathcal{N}$ RGB frames, where $\mathcal{N}\leq K$.
@@ 5.1 Comparison with State-of-the-Art
To evaluate the effectiveness of our proposed architecture, we compare it against several baselines. For pose-only comparisons, we choose a graph-based network MS-G3D [44] and an attention-based method ISTA-Net [66] — the two best performing skeleton-based methods on Assembly101 [52], as reported by [66]. We also employ video baselines TSM [40] and SlowFast [23], emphasizing efficiency and high temporal resolution, respectively. Additionally, we include H2OTR [13], the state-of-the-art for the H2O dataset. For Assembly101, we replicate the results of the methods above using RGB frames from view 4 of the dataset, while results on H2O are acquired from the respective papers. Furthermore, on both datasets, we train and test RGBPoseConv3D [20], which is current state-of-the-art in multimodal action recognition with skeleton and RGB data.
We evaluate three variants of our method by controlling the input modalities. As shown in Tab. 1, our unimodal pose-only model excels in verb recognition on Assembly101. In contrast, RGB-based methods benefit from object appearance and usually perform well for actions due to the strong object recognition.
In this context, the accuracy of ISTA-Net on H2O is not directly comparable to our method, as they also use the 6D object poses, while we only use hand poses as input. RGBPoseConv3D struggles to achieve satisfactory performance, particularly on Assembly101, suggesting that generalizing to hand poses is non-trivial for skeleton-based methods. Therefore, we combine two best-performing unimodal methods with late fusion (MS-G3D + TSM) to create a stronger baseline. Our model even outperforms this, indicating the effectiveness of our proposed multimodal fusion.
Table 1: Quantitative comparison with state-of-the-art methods on Assembly101 [52] and H2O [34]. For H2O, 6D object pose is used by ISTA-Net during training and inference and by H2OTR during training. ‘MS-G3D + TSM’ denotes a late fusion of the corresponding unimodal architectures. Reported GFLOPs include pose estimation costs in the context of Assembly101 at 15 fps. For pose-only methods, pose estimation costs are common and excluded for simpler comparison.
Method Pose RGB GFLOPs Assembly101 H2O Action Verb Object Action MS-G3D [44] 21.2 28.78 63.46 37.26  50.83 ISTA-Net [66] 35.2 28.14 62.70 36.77  89.09 SlowFast [23] 65.7 - - -  77.69 TSM [40] 33.0 35.27 58.27 47.45  - H2OTR[13] - - - -  90.90 RGBPoseConv3D [20] 68.9 33.61 61.99 42.90  83.47 MS-G3D + TSM 66.2 39.74 65.12 51.12  - HandFormer-B/21$\times$8 4.2 28.80 65.33 36.28  57.44 33.0 32.07 55.61 44.89  84.71 47.6 41.06 69.23 51.17 93.39
@@ 5.2 Skeleton-based Action Recognition for Hands
The compositional nature of action classes in hand-object interaction videos allows us to break down the action into a verb and an object. Recognizing such actions from 3D hand poses is an ill-posed problem, as the hand skeletons lack explicit information about the interacting objects, which are also part of the action semantics. However, as the pose data completely captures the hand motion information, it can be reliably used for verb recognition. Therefore, we evaluate a pose-only version of our method on the verb recognition task and compare the performance and efficiency metrics with other state-of-the-art skeleton-based methods in Fig. 5. Our method uses significantly fewer GFLOPs due to our spatiotemporal factorization using micro-actions. With $J=21$, all our HandFormer variants outperform existing methods. The $J=6$ variant is exceptionally efficient with comparable accuracy, notably surpassing the efficiency-focused Shift-GCN [11]. HandFormer-H/21 combines the HandFormer-B and HandFormer-L variants and becomes our best-performing model, improving over MS-G3D by 2.04% while maintaining comparable FLOPs.
@@ 5.3 How many RGB frames are required?
In our model, the pose modality maintains a high temporal resolution to capture fine-grained hand movements, resulting in good verb recognition performance. On the contrary, RGB frames are primarily required to introduce semantic context for object recognition and does not necessitate a high temporal resolution like hand movements. The design of our model allows us to sample only a few RGB frames (as low as one) but still perform competitively at a reduced computational cost.  Fig. 6 shows the impact of using more RGB frames for Assembly101 [52]. For this experiment, non-overlapping micro-actions are considered. Using only one RGB frame in HandFormer (35.46) outperforms the video model TSM (35.27), as shown in Tab. 1. This performance gain stems primarily from a significant improvement in object accuracy, with only a slight enhancement in verb accuracy. However, including more RGB frames shows a diminishing return, which is unsurprising as additional frames are expected to provide redundant information. These results are obtained with a simplified version of our model by setting $\mathcal{L}=\mathcal{L}_{cls}$ and bypassing the multimodal tokenizer.
@@
Figure 5: Comparison of skeleton-based methods for verb recognition on Assembly101 [52]. Our method achieves state-of-the-art performance while utilizing significantly fewer FLOPs.
@@
Figure 6: Ablating number of RGB frames in HandFormer on Assembly101 [52]. With more RGB frames, verb recognition shows marginal gain, whereas object recognition shows improvement with diminishing returns.
5.4 Can 3D hand pose be an efficient alternative to multi-view?
@@
While multiview action recognition benefits from precise hand-movement information in 3D space, processing all views with video models is expensive and highly redundant. In Tab. 2, we demonstrate that combining 3D hand pose with a single RGB view (view 4) achieves comparable performance to multiview action recognition on Assembly101 [52]. Specifically, our action recognition performance (‘Single View + Pose’) matches the fusion of the two most informative views — view 1 and view 4. Notably, our verb recognition accuracy is on par with combining all 8 RGB views, though the latter incurs at least 5×5$\times$5 × more FLOPs despite using the efficient TSM model. Additionally, using hand pose in combination with an RGB view (‘Single View + Pose’) outperforms directly using the egocentric videos (‘Single View + Egocentric’) from which the hand poses are derived. While fusing multiple RGB views improves accuracy by ensembling multiple complementary predictions, the computational overhead increases significantly. In contrast, our model processes hand pose and single-view RGB frames, enhancing efficiency by leveraging the less redundant and low-dimensional pose data.
@@
Table 2: Multi-view Action Recognition on Assembly101[52]. HandFormer with ‘Single View+Pose’ has better performance than ‘Single View+Egocentric’. Also, verb performance is comparable to using all (8) RGB views.
Views Action Verb Object
Single View 35.27 58.27 47.45
Single View + Egocentric 37.75 61.80 49.43
Two Views 41.96 65.22 53.26
All (8) Views 47.51 70.99 57.73
Single View + Pose 41.06 69.23 51.17
Table 3: Cross-view performance of HandFormer in Assembly101 [52] shows its generalization capability to unseen view 1, outperforming the video baseline which is trained on view 1 directly. Egocentric views are the source of hand poses here and, therefore, are included in the video models.
Method Train on Test on Action Verb Object
TSM [40] v4 + ego v4 + ego 37.75 61.80 49.43
v4 + ego v1 + ego 35.27 59.53 47.72
v1 + ego v1 + ego 36.21 60.78 48.52
Our Method v4 + Pose v4 + Pose 41.06 69.23 51.17
v4 + Pose v1 + Pose 38.43 67.86 48.32
@@
Cross-view Generalization. 3D hand pose offers a unique opportunity for cross-view generalization because of its universality across different viewpoints. To evaluate the effectiveness of our method for unseen views, we train our model with frame-wise RGB features from view 4 and test it on view 1. We train TSM [40] as a baseline video model on both RGB views separately. As our method includes 3D hand poses obtained from the egocentric views, we include the egocentric videos in the TSM baselines for a fair comparison. As seen from Tab. 3, our method, trained on view 4, generalizes well on unseen view 1, outperforming the TSM model that was directly trained on view 1.
@@ 5.5 Egocentric Action Recognition
Recognizing actions in egocentric videos is challenging due to camera motion and the occlusion of interacting objects by the hands. Additionally, in the case of Assembly101 [52], the egocentric cameras are similar to Oculus Quest VR headsets, which provide monochrome low-resolution frames. As a result, action recognition performance is significantly lower than fixed RGB views, as found in [52]. We address this challenging scenario with our proposed multimodal architecture and achieve state-of-the-art performance in egocentric action recognition on Assembly101. For this experiment, we used frame-wise TSM features provided by [52]. As depicted in Tab. 4, our model, using a single egocentric view (e3 or e4) outperforms the fusion of four egocentric views as reported in [52]. Moreover, fusing e3 and e4 significantly enhances our model, resulting in a 4.25% increase in action accuracy over the baseline.